# Cuntz–Pimsner algebras of partial automorphisms twisted by vector bundles I: Fixed point algebra, simplicity and the tracial state space

### v2

@@ 2.4. Morphisms and representations of $C^{*}$-correspondences
$\varphi_{\mathcal{F}}(\Pi(a))T(\xi)=T(\varphi_{\mathcal{E}}(a)\xi)\quad\text{for all }a\in A,\xi\in\mathcal{E}.$
@@
Let $(\Pi,T)$ be a morphism of $C^{*}$-correspondences from $\mathcal{E}$ to $\mathcal{F}$. If $\Pi$ is surjective, and if $T$ has dense range and is injective, then $(\Pi,T)$ is covariant.
@@
To finish the proof note that $a\in(\ker\varphi_{\mathcal{E}})^{\perp}$ implies $\Pi(a)\Pi(b)=0$ for all $b\in\ker\varphi_{\mathcal{E}}$. Since $\ker\varphi_{\mathcal{F}}$ is contained in $\Pi(\ker\varphi_{\mathcal{E}})$ this shows that $\Pi(a)c=0$ for all $c\in\ker\varphi_{\mathcal{F}}$, that is, $\Pi(a)\in(\ker\varphi_{\mathcal{F}})^{\perp}$. ∎
@@
A covariant representation of a $C^{*}$-correspondence $\mathcal{E}$ on a $C^{*}$-algebra $B$ is a morphism of $C^{*}$-correspondences from $\mathcal{E}$ to $\mathbf{B}$.
@@
Definition 2.16 ([29, Definition 2.5]).
Let $(\pi,t)$ be a representation of $\mathcal{E}$. We set $t^{0}=\pi$ and $t^{1}=t$. For $n=2,3,...$ we define a linear map $t^{n}:\mathcal{E}^{\otimes n}\to C^{*}(\pi,t)$ by $t^{n}(\xi\otimes\eta)=t(\xi)t^{n-1}(\eta)$ for $x\in\mathcal{E}$ and $\eta\in\mathcal{E}^{\otimes(n-1)}$.
@@ 2.5. Cuntz–Pimsner algebras
The universal covariant representation admits a gauge action which we denote by $\gamma$. The gauge-invariant uniqueness theorem [29, Theorem 6.4], which is fundamental for the theory of Cuntz–Pimsner algebras, says that for a covariant representation $(\pi,t)$ of a $C^{*}$-correspondence $\mathcal{E}$, the $\ast$-homomorphism $\rho:\mathcal{O}(\mathcal{E})\to C^{*}(\pi,t)$ is an isomorphism if and only if $(\pi,t)$ is injective and admits a gauge action. The gauge-invariant uniqueness theorem allows one to show that if $\mathcal{E}$ is a $C^{*}$-subcorrespondence of $\mathcal{F}$, then $\mathcal{O}(\mathcal{E})$ is a subalgebra of $\mathcal{O}(\mathcal{F})$, see [31, Remark 3.7] and [29, Lemma 2.6].
@@
Let $\mathcal{E}$ and $\mathcal{E}_{n}$ for every $n\in\mathbb{N}$ be $C^{*}$-correspondences over a $C^{*}$-algebra $A$. Let $(\Pi_{n},T_{n}):\mathcal{E}_{n}\to\mathcal{E}_{n+1}$ and $(\Pi^{(n)},T^{(n)}):\mathcal{E}^{(n)}\to\mathcal{E}$ be injective covariant morphisms such that $(\Pi^{(n)},T^{(n)})=(\Pi^{(n+1)}\circ\Pi_{n},T^{(n+1)}\circ T_{n})$ for all $n\in\mathbb{N}$. Additionally assume
@@
Since assigning its Cuntz–Pimsner algebra to a $C^{*}$-correspondence gives a covariant functor which preserves monomorphisms (see [40, Proposition 2.11]), we obtain injective *-homomorphisms $\varphi_{n}:\mathcal{O}(\mathcal{E}_{n})\to\mathcal{O}(\mathcal{E}_{n+1})$ and $\varphi^{(n)}:\mathcal{O}(\mathcal{E}_{n})\to\mathcal{O}(\mathcal{E})$ such that $\varphi^{(n)}=\varphi^{(n+1)}\circ\varphi_{n}$ for all $n\in\mathbb{N}$. By the universal property of the inductive limit, this yields a *-homomorphism $\psi:\lim_{\to}\mathcal{O}(\mathcal{E}_{n})\to\mathcal{O}(\mathcal{E})$, which is injective because all $\varphi^{(n)}$ are injective. To show that it is surjective, we need to show
@@ 3.1. Hilbert $C_{0}(X)$-modules
The space $\Gamma_{0}(\mathcal{V})$ with right multiplication $(\xi f)(x)=\xi(x)f(x)$ by continuous functions and the inner product $\langle\xi,\eta\rangle(x)=\beta(\xi(x),\eta(x))$ is a right Hilbert $C_{0}(X)$-module.
@@
Since $(\Phi_{K}(\xi_{k}))_{k}$ converges to $\xi^{(K)}$, there exists $k_{0}>n_{0}$ such that $\left\|\Phi_{K}(\xi_{k})-\xi^{K}\right\|\leq\varepsilon$ for all $k>k_{0}$. Put into words, this means that for all $k>k_{0}$ the restriction of $\xi_{k}$ to $K$ is $\varepsilon$-close to the restriction of $\xi$ to $K$. Since outside of $K$ both $\xi_{k}$ and $\xi$ are bounded by $\varepsilon/2$, this shows $\left\|\xi_{k}-\xi\right\|\leq\varepsilon$. Therefore $(\xi_{n})$ converges to $\xi\in\Gamma_{0}(\mathcal{V})$. ∎
@@
Let $\mathcal{V}$ be a vector bundle over $X$. Recall that we assume the bundle projection to be surjective, and each fiber to be at least one-dimensional. Thus the right Hilbert $C_{0}(X)$-module $\Gamma_{0}(\mathcal{V})$ is full.
@@
We define a $\ast$-homomorphism $R:C_{0}(X)\to\mathcal{L}(\mathcal{E})$ by $R(f)\xi=\xi f$. It is shown in [7, Section 3.2] that $R$ maps into the center of $\mathcal{L}(\mathcal{E})\cong M(\mathcal{K}(\mathcal{E}))$, which gives $\mathcal{K}(\mathcal{E})$ the structure of a $C_{0}(X)$-algebra. The fiber $\mathcal{K}(\mathcal{E})_{x}$ at a point $x\in X$ is isomorphic to $\mathcal{K}(\mathcal{E}_{x})$. The algebra of adjointable operators, being the multiplier algebra of $\mathcal{K}(\mathcal{E})$, fibers over $X$ as well, but it is not a $C_{0}(X)$-algebra (see [3]).
@@
Since the $K_{n}$ exhaust $X$ we can take an approximate unit $(\chi_{n})_{n\in\mathbb{N}}$ of $C_{0}(X)$ such that $\chi_{n}\in C_{0}(\operatorname{int}K_{n})$ for all $n\in\mathbb{N}$. Then $\xi u_{n}\to\xi$ in norm for all $\xi\in\mathcal{E}$. This shows the claim. ∎
@@ 3.2. Partial actions
A set $Y\subset X$ is called $\theta$-invariant if for all $g\in G$ we have $\theta_{g}(Y\cap D_{-1})\subset Y$. A partial action is called minimal if it does not have any closed (equivalently, open) $\theta$-invariant subsets.
For a partial automorphism as in example 3.10, freeness can be expressed in terms of the generating homeomorphism $\theta$. Namely the partial action is free if and only if there is no point $x\in X$ such that $\theta(x)=x$. Invariance of sets can also be expressed in terms of $\theta$, and hence minimality can as well.
@@ 3.3. Partial automorphisms, vector bundles and their $C^{*}$-correspondences
On the other hand, let $f$ be in $C_{0}(X)$ such that $f(x_{0})\neq 0$ for some $x_{0}\in V$. Then there exists an open set $W\subset V$ such that $f(x)\neq 0$ for every $x\in W$. Shrinking $W$ if necessary we may assume there exists an isomorphism $h:\theta^{-1}(W)\times\mathbb{C}^{n}\to\mathcal{V}|_{\theta^{-1}(W)}$ for some $n\in\mathbb{N}$. Let $g\in C_{0}(\theta^{-1}(W))\subset C_{0}(U)$ be nonzero, and let $v$ be any nonzero vector in $\mathbb{C}^{n}$. Then $h(\cdot,g(\cdot)v)$ lies in $\mathcal{E}$, and there is $x\in\theta^{-1}(W)$ such that $\varphi_{\mathcal{E}}(f)h(x,g(x)v))=h(x,g(x)v)f\circ\theta(x)\neq 0$. Hence
@@
The proof is basically the same as the proof of [2, Proposition 3.5]. Let $\{U_{i},h_{i}\}_{i\in I}$ be an an atlas for $\mathcal{V}$. Choose $i,j$ and $k$ in $I$ such that $U_{i}\cap U_{j}\cap U_{k}\neq\emptyset$. Define sections $\xi(x)\coloneqq h_{i}(x,r(x))$, $\eta(x)=h_{j}(x,s(x))$ and $\zeta(x)=h_{k}(x,t(x))$ for continuous functions $r\in C_{0}(U_{i})$, $s\in C_{0}(U_{j})$ and $t\in C_{0}(U_{k})$. Take $x\in U_{i}\cap U_{j}\cap U_{k}$. We have
@@
Assume that there exists $x\in X$ such that $p^{-1}(x)$ has dimension greater then one. There exists a chart $(W,h)$ such that $x$ lies in $W$, where $h:W\times\mathbb{C}^{n}\to p^{-1}(U)$ with $n>1$. Take a matrix $T\in M_{n}(\mathbb{C})$ which is not a multiple of the identity. Let $\tilde{W}$ be an open neighborhood of $x$ contained in $W$, and let $\chi$ be a continuous function which is equal to one on $\tilde{W}$ and vanishes outside of $W$. Let $t$ be the map on $\mathcal{E}$ that maps a section $\xi$ to the section $(t\xi)(y)=h(y,Th^{-1}(\xi(y)))$. One can show that $t$ is well defined and an element of $\mathcal{L}(\mathcal{E})$. Proposition 3.5 shows that $t$ lies in $\mathcal{K}(\mathcal{E})$. However $t$ clearly does not lie in $\varphi_{\mathcal{E}}(C_{0}(V))$.
Now we show the other direction. By Lemma 3.12, Katura’s ideal $J_{\mathcal{E}}$ is equal to $C_{0}(V)$, and hence $\varphi_{\mathcal{E}}(C_{0}(V))$ is contained in $\mathcal{K}(\mathcal{E})$. It is left to show that the restriction of $\varphi_{\mathcal{E}}$ to $J_{\mathcal{E}}$ is surjective onto $\mathcal{K}(\mathcal{E})$. If $\mathcal{V}$ is a line bundle then we have $\xi\langle\eta,\zeta\rangle_{\mathcal{E}}=\zeta\langle\eta,\xi\rangle_{\mathcal{E}}$ for all $\xi,\eta,\zeta\in\mathcal{E}$ by Lemma 3.13. This implies
@@ 3.4. Grading and the fixed point algebra
From the discussion following Definition 2.16, there exists an isometric $\ast$-homomorphism $\psi_{n}$ from $\mathcal{K}(\mathcal{E}^{\otimes})$ to $\mathcal{O}(\mathcal{E})$. On rank-one operators it is given by $\psi_{n}(\theta_{\xi,\eta})=t_{\mathcal{E}}^{n}(\xi)t_{\mathcal{E}}^{n}(\eta)^{*}$ for $\xi,\eta\in\mathcal{E}^{\otimes n}$. For $n\in\mathbb{N}$ we set
@@
Since $\psi_{n}$ is isometric, it is clear that $B_{n}$ is isomorphic to $\mathcal{K}(\mathcal{E}^{\otimes n})$ as $C^{*}$ algebras. Furthermore $B_{[k,n]}$ is an ideal of $B_{[m,n]}$ whenever $m\leq k\leq n$, see [29]. The inductive limit of the increasing sequence $\{B_{[m,n]}\}_{n=m}^{\infty}$ is senoted by $B_{[m,\infty)}$. The $C^{*}$-algebra

### v3

@@ 2.4. Morphisms and representations of $C^{*}$-correspondences
(3) $\varphi_{\mathcal{F}}(\Pi(a))T(\xi)=T(\varphi_{\mathcal{E}}(a)\xi)\quad\text{for all }a\in A,\xi\in\mathcal{E}.$
@@
Let $(\Pi,T)$ be a morphism of $C^{*}$-correspondences from $\mathcal{E}$ to $\mathcal{F}$. If $\Pi$ is surjective and $T$ is bijective, then $(\Pi,T)$ is covariant.
@@
Note that $a\in(\ker\varphi_{\mathcal{E}})^{\perp}$ implies $\Pi(a)\Pi(b)=0$ for all $b\in\ker\varphi_{\mathcal{E}}$. Since $\ker\varphi_{\mathcal{F}}$ is contained in $\Pi(\ker\varphi_{\mathcal{E}})$ this shows that $\Pi(a)c=0$ for all $c\in\ker\varphi_{\mathcal{F}}$, that is, $\Pi(a)\in(\ker\varphi_{\mathcal{F}})^{\perp}$.
We have left to show that $\Pi(\varphi_{\mathcal{E}}^{-1}(\mathcal{K}(\mathcal{E})))$ is contained in $\varphi_{\mathcal{F}}^{-1}(\mathcal{K}(\mathcal{F}))$. But this is clear from surjectivity of $T$ together with Equation 3. It now follows from Definition 2.8 that $\Pi(J_{\mathcal{E}})$ is contained in $J_{\mathcal{F}}$. ∎
@@
Let $\mathcal{E}$ be a $C^{*}$-correspondence over a $C^{*}$-algebra $A$. A covariant representation of $\mathcal{E}$ on a $C^{*}$-algebra $B$ is a covariant morphism of $C^{*}$-correspondences from $\mathcal{E}$ to $\mathbf{B}$.
@@ 2.5. Cuntz–Pimsner algebras
The universal covariant representation admits a gauge action which we denote by $\gamma$. The gauge-invariant uniqueness theorem [29, Theorem 6.4], which is fundamental for the theory of Cuntz–Pimsner algebras, says that for a covariant representation $(\pi,t)$ of a $C^{*}$-correspondence $\mathcal{E}$, the $\ast$-homomorphism $\rho:\mathcal{O}(\mathcal{E})\to C^{*}(\pi,t)$ is an isomorphism if and only if $(\pi,t)$ is injective and admits a gauge action. Here we say that $(\pi,t)$ is injective if $\pi$ is injective, which automatically implies injectivity of $t$. The gauge-invariant uniqueness theorem allows one to show that if $\mathcal{E}$ is a $C^{*}$-subcorrespondence of $\mathcal{F}$, then $\mathcal{O}(\mathcal{E})$ is a subalgebra of $\mathcal{O}(\mathcal{F})$, see [31, Remark 3.7] and [29, Lemma 2.6].
@@
Let $\mathcal{E}$ and $\mathcal{E}_{n}$ for every $n\in\mathbb{N}$ be $C^{*}$-correspondences over $C^{*}$-algebras $A$ and $A_{n}$, respectively. Let $(\Pi_{n},T_{n}):\mathcal{E}_{n}\to\mathcal{E}_{n+1}$ and $(\Pi^{(n)},T^{(n)}):\mathcal{E}_{n}\to\mathcal{E}$ be injective covariant morphisms such that $(\Pi^{(n)},T^{(n)})=(\Pi^{(n+1)}\circ\Pi_{n},T^{(n+1)}\circ T_{n})$ for all $n\in\mathbb{N}$. Additionally assume
@@
Since assigning its Cuntz–Pimsner algebra to a $C^{*}$-correspondence gives a covariant functor which preserves monomorphisms (see [40, Proposition 2.11]), we obtain injective *-homomorphisms $\varphi_{n}:\mathcal{O}(\mathcal{E}_{n})\to\mathcal{O}(\mathcal{E}_{n+1})$ and $\varphi^{(n)}:\mathcal{O}(\mathcal{E}_{n})\to\mathcal{O}(\mathcal{E})$ such that $\varphi^{(n)}=\varphi^{(n+1)}\circ\varphi_{n}$ for all $n\in\mathbb{N}$. By the universal property of the inductive limit, this yields a *-homomorphism $\psi:\lim\limits_{\to}\mathcal{O}(\mathcal{E}_{n})\to\mathcal{O}(\mathcal{E})$, which is injective because all $\varphi^{(n)}$ are injective. To show that it is surjective, we need to show
@@ 3.1. Hilbert $C_{0}(X)$-modules
The space $\Gamma_{0}(\mathcal{V})$ with right multiplication $(\xi f)(x)=\xi(x)f(x)$ by continuous functions vanishing at infinity and the inner product $\langle\xi,\eta\rangle(x)=\beta(\xi(x),\eta(x))$ is a right Hilbert $C_{0}(X)$-module.
@@
Since $(\Phi_{K}(\xi_{k}))_{k}$ converges to $\xi^{K}$, there exists $k_{0}>n_{0}$ such that $\left\|\Phi_{K}(\xi_{k})-\xi^{K}\right\|\leq\varepsilon$ for all $k>k_{0}$. Put into words, this means that for all $k>k_{0}$ the restriction of $\xi_{k}$ to $K$ is $\varepsilon$-close to the restriction of $\xi$ to $K$. Since outside of $K$ both $\xi_{k}$ and $\xi$ are bounded by $\varepsilon/2$, this shows $\left\|\xi_{k}-\xi\right\|\leq\varepsilon$. Therefore $(\xi_{n})$ converges to $\xi\in\Gamma_{0}(\mathcal{V})$. ∎
@@
Let $\mathcal{V}$ be a vector bundle over $X$. Recall that we assume the bundle projection to be surjective, and each fiber to be at least one-dimensional. Using local trivialisations, one can show that each point of $X$ has an open neighborhood $W$ such that $C_{0}(W)$ is contained in $\langle\Gamma_{0}(\mathcal{V}),\Gamma_{0}(\mathcal{V})\rangle$. A partition-of-unity argument implies that $\langle\Gamma_{0}(\mathcal{V}),\Gamma_{0}(\mathcal{V})\rangle$ is equal to $C_{0}(X)$, In other words, $\Gamma_{0}(\mathcal{V})$ is full.
@@
We define a $\ast$-homomorphism $R:C_{0}(X)\to\mathcal{L}(\mathcal{E})$ by $R(f)\xi=\xi f$. It is shown in [7, Section 3.2] that $R$ maps into the center of $\mathcal{L}(\mathcal{E})\cong M(\mathcal{K}(\mathcal{E}))$ and is non-degenerate. By definition, this gives $\mathcal{K}(\mathcal{E})$ the structure of a $C_{0}(X)$-algebra. The fiber $\mathcal{K}(\mathcal{E})_{x}$ at a point $x\in X$ is isomorphic to $\mathcal{K}(\mathcal{E}_{x})$, see [7, Page 72]. The algebra of adjointable operators, being the multiplier algebra of $\mathcal{K}(\mathcal{E})$, fibers over $X$ as well, but it is not a $C_{0}(X)$-algebra (see [3]).
@@
Since the $K_{n}$ exhaust $X$ we can take an approximate unit $(\chi_{n})_{n\in\mathbb{N}}$ of $C_{0}(X)$ such that $\chi_{n}\in C_{0}(\operatorname{int}K_{n})$ for all $n\in\mathbb{N}$. Then $\xi\chi_{n}\to\xi$ in norm for all $\xi\in\mathcal{E}$. This shows the claim. ∎
@@ 3.2. Partial actions
A set $Y\subset X$ is called $\theta$-invariant if for all $g\in G$ we have $\theta_{g}(Y\cap D_{g^{-1}})\subset Y$. A partial action is called minimal if it does not have any closed (equivalently, open) $\theta$-invariant subsets.
A set $Y$ is invariant under a partial automorphism $\theta:U\to V$ as in example 3.10 if and only if we have $\theta(Y\cap U)\subset Y$ and $\theta^{-1}(Y\cap V)\subset Y$.
@@ 3.3. Partial automorphisms, vector bundles and their $C^{*}$-correspondences
On the other hand, let $f$ be in $C_{0}(X)$ such that $f(x_{0})\neq 0$ for some $x_{0}\in V$. Then there exists an open set $W\subset V$ such that $f(x)\neq 0$ for every $x\in W$. Shrinking $W$ if necessary we may assume there exists an isomorphism $h:\theta^{-1}(W)\times\mathbb{C}^{n}\to\mathcal{V}|_{\theta^{-1}(W)}$ for some $n\in\mathbb{N}$. Let $g\in C_{0}(\theta^{-1}(W))\subset C_{0}(U)$ be nonzero, and let $v$ be any nonzero vector in $\mathbb{C}^{n}$. Then $h(\cdot,g(\cdot)v)$ lies in $\mathcal{E}$, and there is $x\in\theta^{-1}(W)$ such that $\varphi_{\mathcal{E}}(f)h(x,g(x)v)=h(x,g(x)v)f\circ\theta(x)\neq 0$. Hence
@@
The proof is basically the same as the proof of [2, Proposition 3.5]. Let $\{U_{i},h_{i}\}_{i\in I}$ be an atlas for $\mathcal{V}$ with transition functions $g_{ij}$. Choose $i,j$ and $k$ in $I$ such that $U_{i}\cap U_{j}\cap U_{k}\neq\emptyset$. Define sections $\xi(x)\coloneqq h_{i}(x,r(x))$, $\eta(x)=h_{j}(x,s(x))$ and $\zeta(x)=h_{k}(x,t(x))$ for continuous functions $r\in C_{0}(U_{i})$, $s\in C_{0}(U_{j})$ and $t\in C_{0}(U_{k})$. Take $x\in U_{i}\cap U_{j}\cap U_{k}$. We have
@@
Assume that there exists $x\in X$ such that $p^{-1}(x)$ has dimension greater than one. There exists a chart $(W,h)$ such that $x$ lies in $W$, where $h:W\times\mathbb{C}^{n}\to p^{-1}(W)$ with $n>1$. Take a matrix $T\in M_{n}(\mathbb{C})$ which is not a multiple of the identity. Let $\tilde{W}$ be an open neighborhood of $x$ contained in $W$, and let $\chi$ be a continuous function which is equal to one on $\tilde{W}$ and vanishes outside of $W$. Let $t$ be the map on $\mathcal{E}$ that maps a section $\xi$ to the section $(t\xi)(y)=h(y,Th^{-1}(\xi(y)))$. One can show that $t$ is well defined and an element of $\mathcal{L}(\mathcal{E})$. Proposition 3.5 shows that $t$ lies in $\mathcal{K}(\mathcal{E})$. However $t$ clearly does not lie in $\varphi_{\mathcal{E}}(C_{0}(V))$.
Now we show the other direction. By Lemma 3.12, Katsura’s ideal $J_{\mathcal{E}}$ is equal to $C_{0}(V)$, and hence $\varphi_{\mathcal{E}}(C_{0}(V))$ is contained in $\mathcal{K}(\mathcal{E})$. It is left to show that the restriction of $\varphi_{\mathcal{E}}$ to $J_{\mathcal{E}}$ is surjective onto $\mathcal{K}(\mathcal{E})$. If $\mathcal{V}$ is a line bundle then we have $\xi\langle\eta,\zeta\rangle_{\mathcal{E}}=\zeta\langle\eta,\xi\rangle_{\mathcal{E}}$ for all $\xi,\eta,\zeta\in\mathcal{E}$ by Lemma 3.13. This implies
@@ 3.4. Grading and the fixed point algebra
Definition 3.19 ([29, Definition 2.5]).
Let $(\pi,t)$ be a representation of $\mathcal{E}$. We set $t^{0}=\pi$ and $t^{1}=t$. For $n=2,3,...$ we define a linear map $t^{n}:\mathcal{E}^{\otimes n}\to C^{*}(\pi,t)$ by $t^{n}(\xi\otimes\eta)=t(\xi)t^{n-1}(\eta)$ for $\xi\in\mathcal{E}$ and $\eta\in\mathcal{E}^{\otimes(n-1)}$.
For $n\in\mathbb{N}$ we set
@@
Since $\psi_{n}$ is isometric, it is clear that $B_{n}$ is isomorphic to $\mathcal{K}(\mathcal{E}^{\otimes n})$ as $C^{*}$ algebras. Furthermore $B_{[k,n]}$ is an ideal of $B_{[m,n]}$ whenever $m\leq k\leq n$, see [29]. The inductive limit of the increasing sequence $\{B_{[m,n]}\}_{n=m}^{\infty}$ is denoted by $B_{[m,\infty)}$. The $C^{*}$-algebra